### v1

Abstract
Climate change exacerbates extreme weather events like heavy rainfall and flooding. As these events cause severe losses of property and lives, accurate high-resolution simulation of precipitation is imperative. However, existing Earth System Models (ESMs) struggle with resolving small-scale dynamics and suffer from biases, especially for extreme events. Traditional statistical bias correction and downscaling methods fall short in improving spatial structure, while recent deep learning methods lack controllability over the output and suffer from unstable training. Here, we propose a novel machine learning framework for simultaneous bias correction and downscaling. We train a generative diffusion model in a supervised way purely on observational data. We map observational and ESM data to a shared embedding space, where both are unbiased towards each other and train a conditional diffusion model to reverse the mapping. Our method can be used to correct any ESM field, as the training is independent of the ESM. Our approach ensures statistical fidelity, preserves large-scale spatial patterns and outperforms existing methods especially regarding extreme events and small-scale spatial features that are crucial for impact assessments.
keywords:
precipitation simulation, statistical downscaling, bias correction, conditional diffusion models, extreme weather events, Earth system modelling
1 Introduction
With global warming, we anticipate more intense rainfall events and associated natural hazards, e.g. in terms of floods and landslides, in many regions of the world [16]. Understanding and accurately simulating precipitation is particularly important for adaptation planning and, hence, for mitigating damages and reducing risks associated with climate change. Earth System Models (ESMs) play a crucial role in simulating precipitation patterns for both historical and future scenarios. However, these simulations are computationally extremely demanding, primarily because they require solving complex partial differential equations. To manage the computational load, ESMs resort to approximate solutions on discretized grids with coarse spatial resolution (typically around 100 km). The consequence is that these models do not resolve small-scale dynamics such as many of the processes relevant to precipitation generation. This leads to considerable biases in ESM fields compared to observations. Moreover, the coarse spatial resolution prevents accurate projections of localized precipitation extremes. The precipitation fields simulated by ESMs can, therefore, not be directly used for impact assessments [28] and especially tasks like water resource and flood management, which demand precise spatial data at high resolution [7].
Statistical bias correction methods can be used as a post-processing to adjust statistical biases. Quantile mapping (QM) is the most common method for improving the statistics of ESM precipitation fields [24, 6, 1]. QM reduces the bias using a mapping that, locally at each grid cell, aligns the estimated cumulative distribution of the model output with the observed precipitation patterns over a reference time period. While QM is effective in correcting distributions of single grid cells, it falls short in improving the spatial structure and patterns of precipitation simulations [10]. Visual inspection shows that ESM precipitation remains too smooth compared to observational data after applying quantile mapping (Fig. 3).
To address these problems, deep learning methods have recently been introduced [11, 18, 4, 10]. In these approaches, the statistical relationships between model simulations and observational data are learned implicitly. A general constraint when using machine learning methods for bias correction is that individual samples of observational and Earth System Model data are always unpaired. A sample, in this context, is a specific weather situation at a specific point in time. The reason for this lack of pairs is that simulations, even with very similar initial conditions, diverge already after a short period of time due to the chaotic nature of the underlying atmospheric dynamics. Currently, one can, therefore, not utilize the wide range of supervised ML techniques that have shown great success in various disciplines in recent years and the available options are consequently restricted to self- and unsupervised machine learning methods. Recent studies [11, 18, 4, 10] applied generative adversarial networks (GANs [5]) and specifically cycleGANs [29] to improve upon existing bias correction techniques. A major limitation of GAN-based approaches is that stability and convergence of the training process are hard to control and that it is challenging to find metrics that indicate training convergence. Also, GANs often suffer from mode collapse, where only a part of the target probability distribution is approximated by the GAN.
As noted above, the low spatial resolution of ESM fields prevents local risk and impact assessment, which necessitates the additional use of downscaling methods. In line with the climate literature, we refer to increasing the spatial resolution as downscaling throughout our manuscript, although we are aware that, especially in the machine learning literature, the term upsampling is more prevalent. We use the term downscaling only when we want to increase the information content in an image as well as the number of pixels. When we refer to upsampling (downsampling), we only mean an increase (decrease) in the number of pixels. Statistical downscaling aims to learn a transformation from the low resolution ESM fields to high resolution observations. Recent developments lean towards using machine learning methods for this task [19, 14]. The potential for machine learning based downscaling methods was shown already in [25]. The work presented a downscaling approach built on purely convolutional networks. More recent work [26] introduced a framework for downscaling and bias correction, combining a diffusion model for downscaling and a model based on optimal transport for bias correction [2]. Optimal transport learns a map between two data distributions in an unsupervised setting. The downside of their framework is its computational expense. The effectiveness also has not been demonstrated on real observational and ESM fields yet.
Diffusion models (DMs) have recently emerged as the state-of-the-art ML approach for conditional image generation [22, 20, 23] and image-to-image translation [21], mostly outperforming GANs across different tasks. Diffusion models (Fig. 1 and S1) avoid the common issues present with GANs in exchange for slower speed at inference. A diffusion model consists of a forward and a backward process. During the forward process, noise is added to an image in subsequent steps to gradually remove its content. The amount of noise added follows a predefined equation. During the backward process, a neural network is trained to reverse each of these individual noising steps to recover the original image. The trained diffusion model can generate an image of the training data distribution, given a noise image as input.
We present here a framework based on state-of-the-art diffusion models that allows to perform both bias correction and downscaling with one single network. This network can, in principle, be any supervised machine learning method. We use a single conditional diffusion model (Fig. 1 and Fig. S1) to correct low resolution (LR) ESM fields towards high resolution (HR) observational data (OBS). In general, our task of bias correction and downscaling can be seen as taking a field from a distribution $p(ESM)$ and transforming it into a field from a conditional distribution $p(OBS|ESM)$. We cannot directly learn such a downscaling and bias correction transformation because ESM and OBS are unpaired. Our framework, therefore, introduces transformations $f$ and $g$ that map ESM and OBS data to a shared embedding space (see Methods and Fig. 1A). On this shared embedding space, we can train a conditional diffusion model to approximate the inverse of $f$. We are then able to bias correct and downscale ESM fields by first mapping them to the shared embedding space and then projecting them to the OBS dataset with the trained diffusion model (Fig. 1B and Fig. 1C). This framework offers great flexibility as it can be applied to any ESM.
Figure 1: (A) Schematic overview of our approach. Bias correction and downscaling can be formulated as a mapping $\omega$ from the $ESM$ data space to the data space of observations ($OBS$) used for training. We first map both datasets to a shared embedding space and then learn the inverse of the mapping $f$ with a diffusion model $DM$. We achieve a correction of the ESM data by applying $DM\circ g$. (B) Our framework allows to train a single model for bias correction and downscaling in a supervised way despite the unpaired nature of observations (OBS) and Earth system model (ESM) fields. We construct functions $f,g$ that map $OBS\in\mathbf{V^{obs}}$ and $ESM\in\mathbf{V^{esm}}$ fields to a shared embedding space $\mathbf{V^{emb}}$. By inverting $f$, we can rewrite $\omega$ as $\omega=f^{-1}\circ g$. We learn the inverse $f^{-1}$ with a conditional diffusion model $DM$. This model is trained (blue arrow) on pairs to approximate the map from $f(OBS)$ to $OBS$ (see Methods for details). Because $f(OBS)$ and $g(ESM)$ share the embedding space (and are identically distributed by construction), we can evaluate (green arrow) the conditional diffusion model $DM$ on the embedded ESM data $g(ESM)$ and thereby approximate the bias correction and downscaling function $\omega=f^{-1}\circ g\approx DM\circ g$.
(C) Left: Training process of the conditional diffusion model $DM\approx f^{-1}$. Note that the individual samples of the input $OBS$ and their embeddings $f(OBS)$, as well as the embeddings $f(OBS)$ and the output of $DM\approx f^{-1}$ are paired, respectively. Right: Inference process of $DM\approx f^{-1}$. In this case, the individual samples of the input $ESM$, their embeddings $g(ESM)$, and the output of $DM\approx f^{-1}$ are paired, respectively. It is not necessary for the training embedding samples to be paired with the inference embedding samples. See Fig. S1 for details.
2 Results
The ability of the diffusion model $DM$ to approximate $f^{-1}$ and the effectiveness of the transformations $f,g$ will determine the overall performance of the downscaling and bias correction model $\omega=DM\circ g$. Therefore, we first investigate the effectiveness of the embedding transformations $f$ and $g$, followed by an analysis of the downscaling and bias correction performance of the diffusion model $DM$, on the observational dataset. Once we have shown that both work as expected, we investigate the performance of the diffusion model in bias correction and downscaling of the ESM precipitation fields. Without loss of generality, we chose the 0.25° ERA5 reanalysis [9] data as observational data and the state-of-the-art GFDL-ESM4 [3] at 1° as our ESM.
2.1 Embedding evaluation
The transformations $f,g$ are chosen such that they map the observational ($OBS$) and model ($ESM$) data to a common embedding space $\mathbf{V^{emb}}$, where all samples are identically distributed. For constructing $f$ and $g$ we need $f(ERA5)$ and $g(GFDL)$ to be unbiased with respect to each other. The transformations need to be chosen such that the embedded data share the same distribution and the same power spectral density (PSD). We assess if they are statistically unbiased towards each other by analyzing their histograms latitude / longitude profiles, as well as their spatial spectra. Figure S2 shows that $f(ERA5)$ and $g(GFDL)$ have the same spatial distribution (Fig. S2A) with minor differences in temporal statistics shown by the histogram (Fig. S2B) and latitudinal/ longitudinal profiles (Fig. S2C and Fig. S2D).
Figure 2: Power spectral densities (PSDs) for different choices noising scale of the diffusion model. The nosing scale $s$ (dashed line) is a hyperparameter that can be chosen depending on the ESM and observational datasets, as well as on the specific task. For the maximal choice of $s$ (A) all information in the observations (ERA5) and model simulations (GFDL) is noised and thereby destroyed. Conditioning on pure noise makes the task equivalent to unconditional image generation. The diffusion model will learn to generate observational fields with no relation to the ESM fields. When $s$ is chosen to be minimal, there will be no noising and the conditional generation will directly replicate the condition, i.e. the ESM field. In (B) we chose $s$ as the point where the PSDs of the observational and simulated datasets intersect. We then apply sufficiently many forward diffusion noising steps to both datasets, destroying small-scale structure until they agree in the PSD. We call scales smaller than $s$ small scales and scales larger than $s$ large scales. In (C) and (D), the effects of choosing a smaller noise scale $s$ are shown. Prior knowledge about the ESM or its accuracy can also guide the choice of $s$.
The individual operations that make up the transformations $f$ and $g$ do not change the large-scale patterns of their respective inputs, as desired for a valid bias correction. The goal of the downscaling and bias correction $\omega$ (Fig. 1) is to rely on the unbiased large-scale patterns of the $ESM$ and correct statistics, as well as small-scale patterns. The transformation $g$ conserves the unbiased information from the ESM by construction. Therefore, we want the diffusion model, approximating $f^{-1}$, to also preserve unbiased information.
The extreme case of erasing all detail with large amounts of noise leads to learning the unconditional distribution $p(ERA5)$, which is then not a correction of $GFDL$ but a generative emulation of the ERA5 reanalysis data. We test this by adding the same amount of noise to the output of our diffusion model that was added to create $g(GFDL)$. This ensures that both the downscaled and bias corrected fields, as well as the original GFDL fields, lack the small-scale details up to the same point.
To verify that large-scale patterns are preserved by the diffusion model, we compute image similarity metrics between the low pass filtered version of the input of the diffusion model (embedded ERA5 data $f(ERA5)$) and the low pass filtered output of the diffusion model $DM(f(ERA5))$. The output of the low pass filter leaves the large-scale features unchanged. The comparison yields an average structural similarity index (SSIM [27]) value of 0.8 and a Pearson correlation coefficient of 0.9 for the validation dataset. This verifies that large-scale patterns are well preserved by the diffusion model.
Our diffusion model is able to reconstruct high-resolution fields following the ERA5 distribution from embedded ERA5 fields $f(ERA5)$, with only minor discrepancies in small-scale patterns (Fig. S2A). A comparison between the mean absolute spatial-temporal difference between GFDL and ERA5 fields yields a bias of 0.3 mm/d. The downscaling and bias correction of our diffusion model reduces this bias to 0.1 mm/d.
Our diffusion model thus approximates $f^{-1}$ well and we successfully created a shared embedding space in which $f(ERA5)$ and $g(GFDL)$ are identically distributed.
2.2 Evaluation of downscaling and bias correction performance
We investigate the inference performance of our diffusion model on embedded GFDL data $g(GFDL)$. We compare the downscaling and bias correction performance of our diffusion model to a benchmark consisting of first applying bi-linear upsampling followed by QM for bias correction.
Figure 3 allows for a qualitative comparison between the different individual precipitation fields. The upsampled $GFDL$ fields, as well as our benchmark consisting of the upsampled and QM-corrected $GFDL$ fields, are visually too smooth. They therefore appear blurry compared to the ERA5 precipitation fields despite being on the same spatial resolution of 0.25°. Our diffusion model produces detailed, high resolution outputs that are visually indistinguishable from the $ERA5$ reanalysis that we treat as ground truth.
Figure 3: Comparative visualization of individual randomly selected samples: Each row presents four samples of the same dataset. The top row shows GFDL ESM4 data, bi-linearly upsampled to 0.25° to match the other fields. The second row shows QM-corrected and the third row diffusion model-corrected GFDL fields. The bottom row shows samples of the original ERA5 data, which are unpaired to the GFDL fields above. Visual inspection shows that the diffusion model correction greatly improves upon the QM correction in terms of producing realistic spatial patterns, since the QM-corrected fields remain way too blurry compared to the HR ERA5 data. There is no visual difference between the details and sharpness of diffusion model-corrected GFDL fields compared to ERA5.
A comparison between the absolute temporally and absolute spatial-temporally averaged diffusion model-corrected GFDL and ERA5 fields (Fig. 4A) yields a bias of 0.29 mm/d. This is a significant improvement over the original GFDL dataset which yields a bias of 0.69 mm/d (Fig. 4B). Our diffusion model performs on par with the state-of-the-art bias correction performance of our benchmark that is by design optimal for this task, at 0.26 mm/d (Fig. 4D).
Figure 4: Bias of the GFDL ESM4 model and the QM- and diffusion model-corrections, defined as the difference between long-term temporal averages of all validation samples. Specifically, the temporally averaged bias fields with respect to ERA5 are shown for (A) the diffusion model correction, (B) the uncorrected GFDL and (C) the QM correction. Results indicate a substantial improvement of our diffusion model (A) and the benchmark (C) over just upsampling GFDL to 0.25°. The absolute bias on top of each panel is given by the mean absolute value of the differences over the spatial and temporal dimension with respect to ERA5.
Figure 5: Comparison of GFDL (bi-linearly upsampled to 0.25°) (orange) and ERA5 (black) to diffusion model-corrected GFDL (magenta) and QM-corrected GFDL fields (blue) as our benchmark. The Power spectral density (PSD) plot (A) shows that the diffusion model corrects the small-scale spatial details far better then our benchmark. The spectrum aligns very well with the high resolution ERA5 target data. The histograms (B) as well as the latitude (C) and longitude (D) profiles show substantial improvements compared to the uncorrected GFDL data.
There are large differences between GFDL and ERA5 data in small-scale patterns (Fig. 5A). The histogram of precipitation intensities (Fig. 5B) also confirms that the ESM is only really accurate for precipitation events up to 40 mm/d, after which the respective frequencies diverge. The latitudinal and longitudinal mean profiles (Fig. 5C and Fig. 5D) indicate the presence of regional biases.
Our framework demonstrates comparable skill to the QM-based benchmark in correcting the latitude and longitude profiles, for which QM is near optimal by construction (Fig. 5C and Fig. 5D). Comparing the histograms (Fig. 5B and Fig. S3) shows a similar performance of our diffusion model compared to the benchmark while even slightly outperforming it for extreme values.
For the spatial patterns and especially the small-scale spatial features, the QM benchmark shows only slight improvements over the original GFDL data (Fig. 5A). The diffusion model is vastly superior in correcting these small-scale spatial patterns (Fig. 5A and Fig. 3) and almost completely removes the small-scale biases, as seen in the spatial PSD.
Our method hence accurately preserves the large-scale precipitation content, while successfully correcting small-scale structure of the precipitation fields, as well as statistical biases in terms of histograms and latitude / longitude profiles (Fig. 5).
3 Discussion
We introduced a framework based on generative machine learning that enables to do both bias correction and downscaling of Earth system model fields with a single diffusion model. We achieve this by first mapping observational fields and ESM data to a shared embedding space and then applying the learned inverse of the observation embedding transformation to the embedded ESM fields. We learn the inverse transformation with a conditional diffusion model. Although the underlying observational and ESM fields are unpaired, our framework allows for training on paired data, and therefore any supervised machine learning method can be adopted to the task, which allows for more flexibility. The diffusion model is trained on individual samples and has successfully learned to reproduce the statistics of observational data. For the observational ground truth, we chose the ERA5 reanalysis, and for the ESM data to be corrected and downscaled, we chose fields from GFDL-ESM4.
Our diffusion model corrects small-scale biases of the ESM fields, while completely preserving the large-scale structures, which is key for impact assessments, especially with regard to extremes and local impacts in terms of floods or landslides. The diffusion model performs particularly well for extreme events where traditional methods struggle. The method improves the temporal precipitation distribution at a grid cell level and surpasses the state-of-the-art approach (quantile mapping) at correcting spatial patterns. The downscaling performance has also been shown to be excellent. The diffusion model manages to generate small-scale details for the low resolution ESM data, that match those of high resolution observations.
Our method is not specific to ERA5 and GFDL, because the training of the diffusion model does not directly depend on the ESM choice. A specific ESM choice will only modify a hyperparameter in the embedding transformations $f$ and $g$. This, however, requires almost no fine-tuning, as the temporal frequencies can always be matched with quantile mapping. The only parameter that might change for different datasets and use cases is the amount of noise that is added to the observational and ESM datasets. We choose the amount of noise such that the PSDs of the observational ground truth and the ESM fields align beyond a certain scale. This means that we have complete flexibility in deciding which patterns we want to preserve and which we want to correct. This is a major advantage over existing GAN based approaches.
We can decrease the level of detail that is preserved by the diffusion model through increasing the amount of noise added in the transformations $f$ and $g$. This leads to higher uncertainties and therefore increased variance in the outputs of the diffusion model. It is, however, not guaranteed that this increase in variance corresponds to the uncertainties of the ESM. The amount of noise added is directly proportional to the freedom the diffusion model has in generating diverse outputs and inversely proportional to the models’ ability to preserve large-scale patterns.
The downscaled and bias corrected fields will automatically inherit time consistency between different samples up to the noising scale. This means that ESM fields showing two successive days will still look like two successive days after the correction.
We focused on precipitation data over the South American continent, because of its heavily tailed distribution and the pronounced spatial intermittency. Especially at small scales, precipitation data is extremely challenging to model and therefore serves as a reasonable choice to show the frameworks capabilities in a particularly difficult setting. Regional data is chosen due to computational constraints, yet the diverse terrain of our study region, encompassing land, sea, and a wide range of altitudes, enables robust testing of the downscaling and bias correction performance, also given the significant biases of the GFDL model in this region. The extension to global scales is straightforward and requires no changes in the architecture. We intend to include more variables in a consistent manner and on a global scale in future research. Optimizing the inference strategy, with speedup techniques like distillation [17], to decrease the sampling time will prove helpful in this context. Taking the growing number of diffusion model variants into consideration, a comparison between different approaches can help establish the best method suited for the general task of bias correcting and downscaling ESM fields.
It is straightforward to extend our methodology to the downscaling and bias correction of numerical or data driven weather predictions on short to medium-range or even seasonal temporal-scales. This would not require any fundamental changes in the architecture. One would, however, require a target dataset with sufficiently high resolution. The ability of the diffusion model to not disturb temporal consistency between samples can be useful in this scenario.
Declarations
Funding
Funded under the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare. This is ClimTip contribution #X; the ClimTip project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101137601. PH, SB, and NB acknowledge funding by the Volkswagen Foundation. BP acknowledges funding by the National Key R&D Program of China (2021YFA0718000). YH acknowledges Alexander von Humboldt Foundation for the Humboldt Research Fellowship.
Data Availability
The ERA5 reanalysis data is available for download at the Copernicus Climate Change Service (https://cds.climate.copernicus.eu/cdsapp#!/dataset/reanalysis-era5-single-levels?tab=overview). The CMIP6 GFDL-ESM4 is available at https://esgf-data.dkrz.de/search/cmip6-dkrz/.
Code Availability
The code will be made available on GitHub.
Competing interests
The authors declare no competing interests.
Ethics approval and consent to participate
Not applicable
Consent for publication
Not applicable
4 Methods
4.1 Data
For the study region, we focus on the South American continent and the surrounding oceans. Specifically, the targeted area spans from latitude 0°N to 63°S and from longitude -90°W to -27°E. The training period comprises ERA5 data from 1992-01-01 to 2011-01-01. The range of years included for evaluation on ERA5 and GFDL spans from 2011-01-02 to 2014-12-01.
ERA5
ERA5 [9] is a state-of-the-art atmospheric reanalysis dataset provided by the European Center for Medium-Range Weather Forecasting (ECMWF). Reanalysis refers to the process of combining observations from various sources, such as weather stations, satellites, and other instruments, with a numerical weather model to create a continuous and comprehensive representation of the Earth’s atmosphere. We use the daily total precipitation data at 0.25° horizontal resolution as the target for the diffusion model.
GFDL
The climate model output is taken from a state-of-the-art ESM from Phase 6 of the Coupled Model Intercomparison Project (CMIP6), namely the GFDL-ESM4 [3]. The dataset contains daily precipitation data of the first ensemble member (r1i1p1f1) of the historical simulation (esm-hist). The data is available from 1850 to 2014, at 1° latitudinal and 1.25° longitudinal resolution and daily temporal resolution.
Benchmark dataset
In order to benchmark our method, we first apply bi-linear interpolation to increase the resolution of the GFDL fields from 1° to 0.25°. After that, we apply Quantile delta mapping [1] to fit the upsampled GFDL data to the original 0.25° ERA5 data. QM is fitted on past observations, and can then be used to correct the statistics of any (past/present) ESM field towards that reference period. We use quantile delta mapping and chose the training period of ERA5 from 1992-01-01 to 2011-01-01 as the reference period for fitting GFDL to ERA5. The benchmark dataset to evaluate our approach is then constructed by applying QM to the validation period of GFDL (2011-01-02 to 2014-12-31).
Data pre-processing
The units of the GFDL data are $\text{kg m}^{-2}\text{s}^{-1}$, and for ERA5 m/h. For consistency, both are transformed to mm/d.
Our pre-processing pipeline consists of:
  • only GFDL: rescaling the original 1°$\times$ 1.25° GFDL data to 1$\times$1° (64$\times$64 pixel).
  • Add +1 mm/d precipitation to each value in order to be able to apply a log-transformation to the data.
  • Apply the logarithm with base 10 in order to compress the range of values.
  • Standardize the data, i.e. subtract the mean and divide by the standard deviation to facilitate training convergence.
  • Transform the data to the range [-1,1] in order to facilitate training convergence.
As part of the transformation $g$, the 1° GFDL data is bi-linearly upsampled. This and the downsampling and upsampling of ERA5 data, which is part of $f$, are already done during pre-processing. The downsampling of 0.25° ERA5 data (256$\times$256 pixel) to 1° (64$\times$64 pixel) is done by only keeping every fourth pixel in each field. For the just mentioned upsampling, we apply bi-linear interpolation to increase the resolution from 1° to 0.25°. Note that bi-linear interpolation to 0.25° does not increase the amount of information in the images compared to the 1° fields. After pre-processing the data as described, the embedding transformation $f$ is applied. The diffusion model is trained with the pre-processed $f(ERA5)$ as a condition and the original 0.25° ERA5 data as a target. Before we apply the embedding transformation $g$ we first pre-process the 1° GFDL data by applying quantile delta mapping (QDM [1]) with 500 quantiles. Bi-linear upsampling is afterward used to increase the resolution to 0.25$\times$0.25° (256$\times$256 pixels). The pre-processed data is used as input to the embedding transformation $g$. The corresponding output serves as the condition during the inference process of the diffusion model.
4.2 Embedding framework
Our framework introduces transformations $f$ $g$ that map OBS and ESM data to a shared embedding space $f:\mathbf{V^{obs}}\rightarrow\mathbf{V^{emb}}$ and $g:\mathbf{V^{esm}}\rightarrow\mathbf{V^{emb}}$. The goal is to do bias correction and downscaling of ESM fields, i.e., to obtain samples from the conditional distribution $\omega=p(OBS|ESM)$. Training a conditional model to approximate this distribution directly is not possible because $OBS$ and $ESM$ are unpaired. The arrows in the diagram of Figure 1 show that we can represent the mapping that achieves the bias correction and downscaling as $\omega=f^{-1}\circ g$. Our idea is to approximate $f^{-1}$ with a neural network $f^{-1}\approx\epsilon$. We chose a conditional diffusion model (DM), denoted by the conditional distribution $p(OBS|f(OBS))$, to approximate $f^{-1}=DM:\mathbf{V^{emb}}\rightarrow\mathbf{V^{obs}}$. The diffusion model (Fig. 1C) is only trained on pairs $(OBS,f(OBS))$. The shared embedding space allows us to evaluate the trained model on ESM embeddings $p(OBS|g(ESM))$, as all embeddings are identically distributed.
4.2.1 Constructing the embedding space
The goal of $f$ and $g$ is to map $OBS$ and $ESM$ to a shared embedding space, where $f(OBS)$ and $g(ESM)$ are identically distributed (Fig. 1). To achieve this, both embedded datasets need to be unbiased towards each other. $OBS$ and $ESM$ are biased towards each other in terms of statistical biases between the distributions and biases between small-scale patterns visible in the spatial power spectral density (PSD) (Fig S2A).
As mentioned earlier, the input to the embedding transformation $f$ is 0.25° ERA5 data, which is first downsampled, then upsampled and pre-processed. The input to the embedding transformation $g$ is the upsampled and pre-processed 0.25° GFDL data. By first downsampling ERA5 to 1° and then upsampling it to 0.25° we ensure that the fields match the information content of the original 1° GFDL fields.
To remove the small-scale pattern bias, we apply a noising procedure analogous to the forward diffusion process as part of $f$ and $g$. Both ERA5 and GFDL data are noised up to a certain spatial scale in the PSD, denoted by the point $s$. The scale $s$ is determined by the point where ERA5 and GFDL start to disagree in the PSD (Fig. 2), i.e., the intersection between the two. Adding noise ensures that $f(ERA5)$ is unbiased compared to $g(GFDL)$ in the PSD by erases all information beneath $s$. The transformations $f$ and $g$ utilize the same cosine scheduler as the forward diffusion process to add Gaussian noise to the data. ERA5 data undergoes 50 noise steps within $f$, while $g$ applies the same 50 noise steps to the GFDL data.
We ensure the observational and ESM data have aligned distributions by incorporating Quantile Mapping (QM) directly into the transformation $g$. It can directly be applied in pre-processing and only needs to be included in $g$. It’s important to clarify that QM is not included because the diffusion model is unable to do bias correction. QM is only used as a tool in our framework to ensure that in the embedding space $f(ERA5)$ and $g(GFDL)$ are identically distributed, such that $g(GFDL)$ can be used for the inference of the diffusion model.
4.2.2 Determining the noising scale
The choice of the spatial scale $s$ influences up to which scale we correct the spatial PSD. We note that the PSD shows spectral distributions normalized to 1; therefore, we can still observe slight changes above $s$ when small-scale patterns are corrected. The point $s$ is a hyperparameter chosen before training and purely depends on the datasets $ESM$ and $OBS$ and can be adjusted to the specific needs in a given context and task.
In the extreme case, where $s$ is maximal, the conditional images will contain pure noise (Fig. 2A). In this case, the diffusion model is equivalent to an unconditional model. As an unconditional model, the diffusion model will correct all biases at all spatial scales, however, at the expense of completely losing any paring between the condition and the output. We chose $s$ to be at the intersection of the ERA5 and GFDL spectrum around 512 km (Fig. 2B). Thereby, we trust in the ESM’s ability to model large-scale structures above the point $s$, which we do not want to correct with the diffusion model.
4.3 Network architecture and training
The general architecture of our diffusion model $DM$ consists of a Denoising Diffusion Probabilistic Model (DDPM) architecture [12] conditioned on low-resolution images. This is similar to a single block of a cascading super-resolution model. We employ current state-of-the-art techniques to facilitate faster convergence and find the following to be important for convergence and sample quality [22]: The memory efficient architecture, ”Efficient U-Net”, in combination with dynamic clipping and noise conditioning augmentation [13] turned out to be effective for our relatively small dataset. We adopt the Min-SNR [8] formulation to weight the loss terms of different timesteps based on the clamped signal-to-noise ratios. The diffusion model architecture utilizes a cosine schedule for noising the target data and a linear schedule for the condition during noise condition augmentation with 1000 steps each. The diffusion model is trained to do v-prediction instead of noise prediction. The U-Net follows the $64\times 64\rightarrow 256\times 256$ Efficient U-Net architecture ([22]). The diffusion model is trained for 80 epochs using the ADAM optimizer [15] with a batch size of 4 and a learning rate of $1e^{-4}$.